OpenStudy (anonymous):

The perimeter of a rectangle is 68 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 9 : 8

OpenStudy (anonymous):

\[2(l+b) = 68\] \[\frac{ l }{ b } = \frac{ 9 }{ 8 }\] \[ l = \frac{ 9 *b }{ 8 }\] Put value of l in first eq